Suppose that a couple in planning 3 children is primarily interested in the number of boys.

1) - Define the random variable X.

2) - What is the range of X?

3) - Find the probability function of X.

4) - Find the probability mass function

5) - Draw the graph for the probability mass function




1
Expert's answer
2022-05-26T08:09:38-0400

1) Let random variable X be  number of boys.


2)


"S=\{BBB, BBG, BGB, GBB,"

"BGG,GBG, BGG, GGG\}"


"Range(X)=\{0,1,2,3\}"

3)


"P(X=0)=\dfrac{1}{8}"

"P(X=1)=\dfrac{3}{8}"

"P(X=2)=\dfrac{3}{8}"

"P(X=3)=\dfrac{1}{8}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 & 3 \\ \hline
 p(x) & 1/8 & 3/8 & 3/8 & 1/8 \\
\end{array}"

4)


"f(x) = \begin{cases}
 1/8, &x=0 \\
 3/8, &x=1 \\
 3/8, &x=2 \\
 1/8, &x=3 \\
\end{cases}"

5)